How many excess electrons must be added to a plastic solid sphere that has a diameter of 0.26 m to produce an electric field of 1030 N/C at the edge of the surface? Assuming the charge is uniform, what is the electric field at 10 cm from the center?



1
Expert's answer
2021-10-14T11:32:39-0400

Find the charge of the solid sphere of charged uniformly:


"Q=4\pi\epsilon_0R^2E."

The number of electrons:


"N=\frac Qe=\frac{4\pi\epsilon_0R^2E}{e}=12·10^9."

The field 10 cm from the center of the sphere:


"E_i=Qr/(4\pi\epsilon_0R^3)=794\text{ N/C}."